### v4

# Pragmatic lossless compression: Fundamental limits and universality

Abstract
The problem of variable-rate lossless data compression is considered, for codes with and without prefix constraints. Sharp bounds are derived for the best achievable compression rate of memoryless sources, when the excess-rate probability is required to be exponentially small in the blocklength. Accurate nonasymptotic expansions with explicit constants are obtained for the optimal rate, using tools from large deviations and Gaussian approximation. When the source distribution is unknown, a universal achievability result is obtained with an explicit “price for universality” term. This is based on a fine combinatorial estimate on the number of sequences with small empirical entropy, which might be of independent interest. Examples are shown indicating that, in the small excess-rate-probability regime, the approximation to the fundamental limit of the compression rate suggested by these bounds is significantly more accurate than the approximations provided by either normal approximation or error exponents. The new bounds reinforce the crucial operational conclusion that, in applications where the blocklength is relatively short and where stringent guarantees are required on the rate, the best achievable rate is no longer close to the entropy. Rather, it is an appropriate, more pragmatic rate, determined via the inverse error exponent function and the blocklength.
00footnotetext: All three authors are with the Statistical Laboratory, Centre for Mathematical Sciences, University of Cambridge, Wilberforce Road, Cambridge CB3 0WB, UK. Email: at771@cam.ac.uk, lg560@cam.ac.uk, yiannis@maths.cam.ac.uk.
This work was supported in part by the EPSRC-funded INFORMED-AI project EP/Y028732/1.
A preliminary version of some of the results in this paper appeared at ISIT 2025 [36].
Keywords — Data compression, memoryless source, error exponent, normal approximation, large deviations, pragmatic rates, universal compression, types, empirical entropy
1 Introduction
1.1 Variable-rate data compression
Let $\mbox{$X$}=\{X_{n}\;;\;n\geq 1\}$ be a memoryless source consisting of independent and identically distributed (i.i.d.) random variables $X_{n}$, with values in a finite alphabet $A$, and with common distribution described by the probability mass function (p.m.f.) $P$: ${\mathbb{P}}(X_{n}=x)=P(x)$, $x\in A$.
We revisit the problem of losslessly compressing the output of such a memoryless source, when its distribution is known a priori as well as when it is not. The problem formulation is simple and quite elementary, though its importance can hardly be overstated, in view of its application across the sciences and engineering.
For strings of symbols from $A$, we write $x^{n}$ for the block $x^{n}=(x_{1},x_{2},\ldots,x_{n})$, and similarly, for blocks of random variables, $X^{n}$ denotes $(X_{1},X_{2},\ldots,X_{n})$. A variable-rate code, or fixed-to-variable compressor, for strings of length $n$ from $A$ is an injective function $f_{n}:A^{n}\to\{0,1\}^{*}$, where $\{0,1\}^{*}$ denotes the set of all finite-length binary strings:
$\{0,1\}^{*}=\{\varnothing,0,1,00,01,10,11,000,001,\ldots\}.$
Writing $\ell(s)$ for the length of a string $s\in\{0,1\}^{*}$, the compressor $f_{n}$ maps strings $x^{n}\in A^{n}$ into binary strings $f_{n}(x^{n})$ of length $\ell(f_{n}(x^{n}))$ bits. We call such compressors one-to-one codes.
[…]
$\epsilon_{n}^{*}(R,P):=\min_{f_{n}}\mathbb{P}\big(\ell(f_{n}(X^{n}))\geq nR\big).$
This minimum is achieved by an optimal compressor $f_{n}^{*}$ independent of the rate $R$ [20]. Similarly, for a blocklength $n\geq 1$ and excess-rate probability $\epsilon>0$, the best achievable rate $R_{n}^{*}(\epsilon,P)$, is:
$R_{n}^{*}(\epsilon,P):=\inf\big\{R>0:\min_{f_{n}}\mathbb{P}\big(\ell(f_{n}(X^{n}))\geq nR\big)\leq\epsilon\big\}.$
Prefix-free codes. In the case of prefix-free codes, analogous fundamental limits, $\epsilon^{\sf p}_{n}(R,P)$ and $R_{n}^{\sf p}(\epsilon,P)$, can be defined, by restricting the corresponding minimisations to codes that satisfy the prefix condition. But $R_{n}^{\sf p}(\epsilon,P)$ is very tightly coupled with $R_{n}^{*}(\epsilon,P)$: As shown in [20, Theorem 1], we always have $R^{\sf p}_{n}(\epsilon,P)=R_{n}^{*}(\epsilon,P)+\frac{1}{n}$.
1.2 Background
When the source distribution $P$ is known, and the excess-rate probability $\epsilon$ is fixed, the problem of determining the first-order asymptotic behaviour of the optimal rate $R_{n}^{*}(\epsilon,P)$ essentially goes back to Shannon. We have,
$nR_{n}^{*}(\epsilon,P)=nH(P)+o(n)\;\;\mbox{bits},\quad\mbox{as}\;n\to\infty,$ (1)
where $H(P)=-\sum_{x}P(x)\log P(x)$ as usual denotes the entropy, and, throughout, ‘log’ denotes $\log_{2}$, the logarithm taken to base 2; see [20] for an extensive discussion and historical comments.
The first-order relation (1) was refined by Strassen [33] who claimed that, as $n\to\infty$,
$nR_{n}^{*}(\epsilon,P)=nH(P)+\sqrt{n}\sigma(P)Q^{-1}(\epsilon)-\frac{1}{2}\log n+O(1)\;\;\mbox{bits},$ (2)
where $\sigma^{2}(P)=\mbox{\rm Var}_{P}(-\log P(X))$ is the minimal coding variance [19] or source dispersion [20], and $Q(x)=1-\Phi(x)$, $x\in{\mathbb{R}}$, denotes the standard normal tail function. In fact, Strassen identified the $O(1)$ term explicitly, giving an expansion with an $o(1)$ error term. Although some issues of rigour were raised in [20] regarding his proof, a stronger version of (2) was established in [20], where explicit, finite-$n$ bounds were obtained for the $O(1)$ term.
When the excess-rate probability is not fixed but is instead required to decay to zero exponentially fast, the best rate that can be achieved turns out to be higher than the entropy. Specifically, for $\delta>0$ in an appropriate range, as $n\to\infty$,
$nR_{n}^{*}(2^{-n\delta},P)=nH(P_{\alpha^{*}})+o(n)\;\;\mbox{bits},$ (3)
for a specific $\alpha^{*}\in(0,1)$ depending on $\delta$, where, for each $\alpha\in(0,1)$, $P_{\alpha}$ denotes the tilted distribution,
$P_{\alpha}(x)=\frac{P(x)^{\alpha}}{\sum_{y\in A}P(y)^{\alpha}},\quad x\in A.$
The first-order asymptotic relation (3) is simply the “rate version” of the well-known error-exponents result that was established in the early works [13, 18] and in the present form by Blahut [7]. Indeed, $H(P_{\alpha^{*}})$ can be expressed as the inverse of the error-exponent function,
$\Delta_{P}(R):=\inf_{P^{\prime}:H(P^{\prime})\geq R}D(P^{\prime}\|P),$
via
$H(P_{\alpha^{*}})=\Delta_{P}^{-1}(\delta)=\sup_{P^{\prime}:D(P^{\prime}\|P)\leq\delta}H(P^{\prime}),$
where $D(P^{\prime}\|P)$ is the relative entropy between $P^{\prime}$ and $P$.
1.3 Compression at pragmatic rates: Known source
The first main contribution of the present work is a nonasymptotic refinement of the evaluation of the fundamental compression limit $R_{n}^{*}$, in the small excess-rate probability regime, when the source distribution $P$ is known and fixed. In the same sense in which Strassen’s expansion (2) and the finite-$n$ bounds of [20] strengthen the classical Shannon asymptotic relation (1), we show it is possible to derive finite-$n$ bounds that provide a correspondingly finer and stronger version of the first-order asymptotic expansion (3) in the small excess-rate probability regime.
In Theorems 3.1 and 3.2 of Section 3 we provide explicit constants $C$ and $C^{\prime}$ such that,
$C^{\prime}\leq nR_{n}^{*}(2^{-n\delta},P)-\Big[nH(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\log n\Big]\leq C,$ (4)
for all $n$ greater than some explicit $N_{0}$. In view of this, we may call the expression
${\cal R}_{n}(\epsilon,P):=H(P_{\alpha^{*}})-\frac{1}{2n(1-\alpha^{*})}\log n\;\;\mbox{bits/symbol},$
the best pragmatic rate that can be achieved at blocklength $n$ with excess-rate probability no greater than $\epsilon=2^{-n\delta}$.
On account of the remark at the end of Section 1.1, the exact same bounds as in (4) hold in the case of codes with prefix constraints – that is, for $R_{n}^{\sf p}$ in place of $R_{n}^{*}$ – with the same constant $C^{\prime}$, and with $C+1$ in place of $C$.
Despite their technical nature, the approximations to $R_{n}^{*}$ and $R_{n}^{\sf p}$ provided by all these different approaches are of as much practical relevance as they are of mathematical interest. In particular, they each are useful in different regimes of the blocklength and the excess-rate probability requirements. For example, the Strassen-style approximation suggested by (2) is only relevant for moderate values of $\epsilon$. If $\epsilon$ is small, then the second term dominates and the approximation is no longer valid or useful.
In that case, the operational utility of (4) can be described as follows. Given the blocklength $n$ and a small target excess-rate probability $\epsilon$, we can compute $\delta=(1/n)\log(1/\epsilon)$ and solve for the corresponding $\alpha^{*}$. Then the best achievable rate with these parameters is approximately ${\cal R}_{n}(\epsilon,P)$ bits/symbol. As a simple, concrete example, consider the case of binary memoryless source $\mbox{$X$}=\{X_{n}\}$ where each $X_{n}$ has p.m.f. $P\sim\mbox{\rm Bern}(0.2)$.
We examine the four approximations to $R_{n}^{*}(\epsilon,P)$ suggested by the above results for various small values of $\epsilon$. In the ‘Shannon regime’,
Shannon: $R_{n}^{*}\approx H(P),$ (5)
Strassen: $R_{n}^{*}\approx H(P)+\frac{1}{\sqrt{n}}\sigma(P)Q^{-1}(\epsilon)-\frac{1}{2n}\log n,$ (6)
and in the ‘error exponents regime’,
Blahut: $R_{n}^{*}\approx H(P_{\alpha^{*}}),$ (7)
Thms. 3.1-3.2: $R_{n}^{*}\approx{\cal R}_{n}=H(P_{\alpha^{*}})-\frac{1}{2n(1-\alpha^{*})}\log n.$ (8)
Table 1 shows representative results when the blocklength $n=50$, which clearly demonstrate that the pragmatic rate ${\cal R}_{n}(\epsilon,P)$ provides by far the most accurate estimate of the optimal rate $R_{n}^{*}(\epsilon,P)$ for the problem parameters considered.
$\epsilon$ $R_{n}^{*}(\epsilon,P)$ Shannon (5) Strassen (6) Blahut (7) ${\cal R}_{n}(\epsilon,P)$ (8)
0.00003 0.940 0.722 1.119 1.000 0.941
0.00010 0.940 0.722 1.086 0.997 0.936
0.00032 0.920 0.722 1.052 0.993 0.928
0.00093 0.900 0.722 1.017 0.987 0.917
0.00251 0.900 0.722 0.983 0.979 0.903
0.00626 0.880 0.722 0.948 0.969 0.888
0.01444 0.840 0.722 0.913 0.957 0.869
Table 1: Comparison between the true value of the optimal rate $R_{n}^{*}(\epsilon,P)$ and four different approximations. Clearly the approximation suggested by Theorems 3.1 and 3.2, given by the pragmatic rate ${\cal R}_{n}$, gives the best results in this regime.
1.4 Universal compression at pragmatic rates
[…]
In the Shannon regime, Kosut and Sankar [22] showed that the same optimal performance as for known sources [33, 20] could be achieved universally by one-to-one codes, at a cost of approximately
$\Big(\frac{m-2}{2}\Big)\log n\quad\mbox{bits},$
over all memoryless sources on an alphabet of size $m=|A|\geq 2$. In Theorem 5.1 in Section 5, we prove an analogous result in the error exponents regime. Namely, we describe a specific sequence of one-to-one compressors $\{\phi_{n}^{*}\}$ that simultaneously achieve excess-rate probability no greater than $2^{-n\delta}$ on any memoryless source with full support on $A$, at a rate no greater than:
$H(P_{\alpha^{*}})+\Big(\frac{m-2}{2}-\frac{1}{2(1-\alpha^{*})}\Big)\frac{\log n}{n}+O\Big(\frac{1}{n}\Big)\quad\mbox{bits per symbol}.$ (9)
Interestingly, the code used to prove Theorem 5.1 is somewhat simpler than the compressors employed in [22]. The idea is to simply order all string $x^{n}\in A^{n}$ in order of increasing empirical entropy, and take $\phi_{n}^{*}(x^{n})$ to be the binary description of the index of $x^{n}$ in this ordering.
On the other hand, in the present setting the analysis of their performance requires a very accurate combinatorial estimate on the number of strings with low empirical entropy. Specifically, in Theorem 4.1 we show that the number of $x^{n}\in A^{n}$ such that $H(\hat{P}_{x^{n}})\leq H(Q)$ is:
$\Theta\big(n^{\frac{m-3}{2}}2^{nH(Q)}\big).$
The proof of this estimate is the most technical part of this paper, and the result itself may be of independent interest.
Although we do not provide a matching converse to the achievability statement in Theorem 5.1, we conjecture that the rate obtained there, given in (9), is optimal.
At first glance, it may be surprising that the cost of universality, both in the Shannon regime and in the error exponents regime, is $(\frac{m-2}{2})\log n$ bits, instead of the more familiar $(\frac{m-1}{2})\log n$ bits seen in the classical universal compression literature on compression redundancy. This discrepancy comes from the fact that, unlike, e.g., in Shtarkov’s [32] or Rissanen’s [30] well-known results, we consider one-to-one codes instead of prefix-free codes. In the case of a known source, as describer earlier, this makes a difference of at most one bit [20]. But in the case of universal compression, the difference between optimal prefix-free and optimal one-to-one codes has been found to be of the order of $\log n$ bits, both in the Shannon excess-rate regime [22], and in terms of redundancy rates [6].
One way to see why there is, indeed, a difference in universal coding with one-to-one versus with prefix-free codes, is to consider binary memoryless sources. As described in [20], the optimal one-to-one code of any such source simply orders all strings of length $n$ in decreasing probability, and assigns binary descriptions of increasing lengths to their indices lexicographically. But for binary ${\rm Bern}(p)$ sources there are only two optimal orderings, depending on whether $p$ is greater or smaller than 1/2. Therefore, a universal one-to-one code only needs one extra bit to differentiate between these to cases. On the other hand, the optimal prefix-free universal code needs to describe the maximum likelihood estimate of $p$ based on $x^{n}$, to $\frac{1}{\sqrt{n}}$ accuracy, which takes about $\frac{1}{2}\log n$ bits; see, e.g., the discussions in [4] or [15]. So there is an $O(\log n)$ universality penalty for prefix-free codes, whereas there is only an $O(1)$ universality penalty for one-to-one codes.
We do not consider the problem of prefix-free universal codes in the error exponents regime in this work, but we conjecture that, compared to the optimal rate in (4), they incur a penalty of, at best, $\frac{m}{2}\log n$ bits; cf. [22].
1.5 Prior related work
Formally, the problem of lossless data compression is very closely related to binary hypothesis testing, especially in the case of fixed-rate compression. Indeed, there are close hypothesis-testing parallels to all the major data-compression steps outlined in Section 1.2, beginning with Stein’s lemma [8, 23, 9], its refinement by Strassen [33] and Polyanskiy-Poor-Verdú [29], and the error-exponent asymptotics of Hoeffding [17].
For hypothesis testing, sharper expansions in the error-exponents regime have been developed by numerous authors, including the work reported in [11, 1, 2, 37, 26, 27]. The results in this paper can be viewed as lossless compression analogues of the refined error-exponent bounds obtained for binary hypothesis testing in [26, 27, 25, 24].
For fixed-rate data compression, Csiszár and Longo [11] obtained an asymptotic expansion for the excess-rate probability which is accurate up to and including the $O(1)$ term, when the target rate is fixed. B. Nakiboğlu [26] observed that this $O(1)$ term is incorrect in [11], and there are also further questions of rigour regarding the proofs there, partly stemming from the fact that they rely on earlier results of Strassen [33]. Our Theorems 3.1 and 3.2 can be viewed as finite-$n$ improvements of the “rate version” of these expansions, in the variable-rate case. Also, in a related but different direction, the best achievable rate in the “moderate deviations” regime was examined in [3].
Extensive discussions of different aspects of the fundamental limits of lossless data compression can be found in Csiszár and Körner’s classic text, [10], Han’s book on information spectrum methods [16], and Tan’s monograph [34]. Universal lossless data compression also has a rich history. The texts [31, 12, 15, 9] and the papers [30, 5] provide comprehensive overviews.
2 Preliminaries and auxiliary results
Throughout, $\mbox{$X$}=\{X_{n}\}$ denotes a memoryless source with distribution $P$ that has full support on a finite alphabet $A$. For $\alpha\in(0,1)$, the tilted distribution $P_{\alpha}$ is given by
$P_{\alpha}(x)=\frac{P(x)^{\alpha}}{Z_{\alpha}},\quad x\in A,$ (10)
where the normalising constant $Z_{\alpha}=\sum_{x\in A}P(x)^{\alpha}$. The standard normal cumulative distribution function is denoted by $\Phi$, $\log$’ denotes the logarithm taken to base 2, and the relative entropy between two p.m.f.s $P,Q$ on the same finite alphabet $A$ is $D(P\|Q)=\sum_{x\in A}P(x)\log[P(x)/Q(x)].$
2.1 Normal approximation, compression and hypothesis testing
Here we collect some classical normal approximation inequalities, a simple one-shot converse bound for data compression and some useful nonasymptotic results for binary hypothesis testing, all of which will be employed in the proofs of our results in Sections 3 and 5.
First, we recall the classical Berry-Esséen bound [21, 28] for sums of i.i.d. random variables.
Theorem 2.1 (Berry-Esséen bound)
Let $Z_{1},Z_{2},\ldots,Z_{n}$ be i.i.d. random variables with mean $\mu=\mathbb{E}(Z_{1})$, variance $\sigma^{2}=\textup{Var}(Z_{1})$ and $\rho=\mathbb{E}(|Z_{1}-\mu|^{3})<\infty$. Then:
$\Bigg|\mathbb{P}\Bigg(\frac{\sqrt{n}}{\sigma}\Big(\frac{1}{n}\sum_{i=1}^{n}Z_{i}-\mu\Big)\leq x\Bigg)-\Phi(x)\Bigg|\leq\frac{\rho}{2\sigma^{3}\sqrt{n}},\hskip 14.45377ptx\in\mathbb{R},\hskip 5.05942ptn\geq 1.$
The following is a simple corollary of Theorem 2.1, cf. [29, 25, 24].
Lemma 2.2
Let $Z_{1},Z_{2},\ldots,Z_{n}$ be as in Theorem 2.1. Then:
$\mathbb{E}\Bigg\{\exp\Big(-\sum_{i=1}^{n}Z_{i}\Big)\mathbb{I}_{\{\sum_{i=1}^{n}Z_{i}\geq x\}}\Bigg\}\leq\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho}{\sigma^{2}}\Big)\frac{1}{\sigma\sqrt{n}}e^{-x},\hskip 14.45377ptx\in\mathbb{R},\hskip 5.05942ptn\geq 1.$
The following one-shot converse bound for data compression was established in [20].
Theorem 2.3 (One-shot converse)
Let $f^{*}$ be the optimal compressor for a finite-valued random variable $X$. Then, for any $k>0$ and $\tau>0$,
$\mathbb{P}[\ell(f^{*}(X))\geq k-1]>\mathbb{P}[-\log P(X)\geq k+\tau]-2^{-\tau}.$
The next two results were originally developed in the context of binary hypothesis testing. Lemma 2.4 [11] can be viewed as a “change of measure” result for hypothesis testing, and Theorem 2.5 [25, 24, 29] is a strong, finite-$n$ converse bound for hypothesis testing.
Lemma 2.4 (Change of measure)
Let $P,Q$ be two p.m.f.s with full support on $A$, and let $S_{n}$ be an arbitrary subset of $A^{n}$, for some $n\geq 1$. If
$\frac{P^{n}(x^{n})}{Q^{n}(x^{n})}\geq\max_{y^{n}\in S_{n}}\frac{P^{n}(y^{n})}{Q^{n}(y^{n})}\quad\mbox{for all}\;x^{n}\not\in S_{n},$
then
$P^{n}(S_{n})=\min_{S\subset A^{n}:Q^{n}(S^{c})\leq Q^{n}(S_{n}^{c})}P^{n}(S).$
Theorem 2.5 (Hypothesis testing converse)
Let $P\neq Q$ be p.m.f.s with full support on $A$, and for $n\geq 1$, $\epsilon\in(0,1)$, let
$e_{1}^{*}(\epsilon)=\min_{S\subset A^{n}:Q^{n}(S^{c})\leq\epsilon}P^{n}(S).$
Then, for any $\Delta>0$ such that $\epsilon+\frac{B+\Delta}{\sqrt{n}}<1$,
$\log e_{1}^{*}(\epsilon)\geq-nD(Q\|P)-\sigma\sqrt{n}\Phi^{-1}\Big(\epsilon+\frac{B+\Delta}{\sqrt{n}}\Big)-\frac{1}{2}\log n+\log\Delta,$
where $\sigma^{2}=\textup{Var}_{Q}(\log\frac{Q(X)}{P(X)})$, $\rho=\mathbb{E}_{Q}|\log\frac{Q(X)}{P(X)}-\mathbb{E}_{Q}(\log\frac{Q(X)}{P(X)})|^{3}$ and $B=\frac{\rho}{2\sigma^{3}}$.
2.2 Entropy, moments and derivatives identities
We close with some simple technical lemmas, stated without proof. Let $X\sim P$ where $P$ has full support on $A$, and recall the definition of $P_{\alpha}$ in (10). The proofs of Lemmas 2.62.9 follow by direct computation; see for example, [11] and [25, 24].
Lemma 2.6
Define the tilted second and third moments as:
$\sigma_{2,\alpha}^{2}$ $=\textup{Var}_{P_{\alpha}}\Big(\log_{e}\frac{P_{\alpha}(X)}{P(X)}\Big),$
$\rho_{2,\alpha}$ $=\mathbb{E}_{P_{\alpha}}\Big|\log_{e}\frac{P_{\alpha}(X)}{P(X)}-\mathbb{E}_{P_{\alpha}}\Big(\log_{e}\frac{P_{\alpha}(X)}{P(X)}\Big)\Big|^{3}\,$,
$\sigma_{3,\alpha}^{2}$ $=\textup{Var}_{P_{\alpha}}(\log_{e}P(X)),$
$\rho_{3,\alpha}$ $=\mathbb{E}_{P_{\alpha}}|\log_{e}P(X)-\mathbb{E}_{P_{\alpha}}(\log_{e}P(X))|^{3}.$
Then, for all $\alpha\in(0,1)$: $\sigma_{2,\alpha}^{2}=(1-\alpha)^{2}\sigma_{3,\alpha}^{2}$ and $\rho_{2,\alpha}=(1-\alpha)^{3}\rho_{3,\alpha}$.
Lemma 2.7
For all $\alpha\in(0,1)$ we have the following expressions for the derivatives of $D(P_{\alpha}\|P)$ and $\sigma^{2}_{3,\alpha}$:
$\frac{dD(P_{\alpha}\|P)}{d\alpha}$ $=(\alpha-1)\sigma_{3,\alpha}^{2}(\log e),$
$\frac{d^{2}D(P_{\alpha}\|P)}{d\alpha^{2}}$ $=(\log e)\sigma_{3,\alpha}^{2}+(\log e)(\alpha-1)\frac{d\sigma_{3,\alpha}^{2}}{d\alpha},$
$\frac{d\sigma_{3,\alpha}^{2}}{d\alpha}$ $=\mathbb{E}_{P_{\alpha}}\Big[\Big(\log_{e}P(X)-\mathbb{E}_{P_{\alpha}}[\log_{e}P(X)]\Big)^{3}\Big].$
Lemma 2.8
$D(P_{\alpha}\|P)$ is a continuous, strictly decreasing function of $\alpha$, for $\alpha\in(0,1)$.
Lemma 2.9
For all $\alpha\in(0,1)$ we have the following expressions for the derivatives of $H(P_{\alpha})$:
$\frac{dH(P_{\alpha})}{d\alpha}$ $=-(\log e)\alpha\sigma_{3,\alpha}^{2},$
$\frac{d^{2}H(P_{\alpha})}{d\alpha^{2}}$ $=-(\log e)\Big[\sigma_{3,\alpha}^{2}+\alpha\frac{d\sigma_{3,\alpha}^{2}}{d\alpha}\Big].$
3 Compression of a known source
We will now proceed with the first set of our main results of this paper, namely the development of non-asymptotic achievability and converse bounds on the optimal compression rate, $R_{n}^{*}(2^{-n\delta},P)$, when the excess-rate probability decays exponentially fast.
3.1 Achievability
Theorem 3.1 (Achievability)
[…]
$R_{n}^{*}(2^{-n\delta},P)\leq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}+\frac{c}{n},$ (11)
where $P_{\alpha}$ is defined in (10), $\alpha^{*}$ is the unique $\alpha\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$, and
$c=\log\Bigg(\frac{1}{\sigma_{1}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1}}{\sigma_{1}^{2}}\Big)\Bigg)+\frac{\alpha^{*}}{1-\alpha^{*}}\log\Bigg(\frac{1}{\sigma_{2}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2}}{\sigma_{2}^{2}}\Big)\Bigg),$
where $\sigma_{1}^{2}=\textup{Var}_{P_{\alpha^{*}}}(\log_{e}P_{\alpha^{*}}(X))$, $\rho_{1}=\mathbb{E}_{P_{\alpha^{*}}}|\log_{e}P_{\alpha^{*}}(X)-\mathbb{E}_{P_{\alpha^{*}}}(\log_{e}P_{\alpha^{*}}(X))|^{3}$, $\sigma_{2}^{2}=\sigma_{2,\alpha^{*}}^{2}$ and $\rho_{2}=\rho_{2,\alpha^{*}}$, with $\sigma_{2,\alpha}^{2}$ and $\rho_{2,\alpha}$ defined as in Lemma 2.6.
Remark. As noted in the introduction, the exact bound (11) remains valid in the case of codes with prefix constraints – i.e., for $R_{n}^{\sf p}$ in place of $R_{n}^{*}$ – with $c+1$ in place of $c$.
Proof. The proof is based on an explicit construction of a compressor $f_{n}$. To that end, let $0<\beta_{n}<1$ be a constant that will be chosen later, and define the set
$E_{n}=\{x^{n}\in A^{n}:\log P^{n}(x^{n})\geq\log\beta_{n}\}.$ (12)
Write
$M_{X}(\beta_{n}):=|E_{n}|,$
for the cardinality of $E_{n}$, and suppose that
$P^{n}(E_{n}^{c})\leq 2^{-n\delta}.$ (13)
To obtain the desired compressor on $A^{n}$, first we order the strings $x^{n}\in E_{n}$ in order of decreasing probability $P^{n}(x^{n})$, breaking ties arbitrarily, and define $f_{n}$ via:
[…]
Note that the longest codeword $f_{n}(x^{n})$ among all strings in $E_{n}$ has length $\lfloor\log M_{X}(\beta_{n})\rfloor$. Therefore, for any fixed $\xi\in(0,1)$,
$\mathbb{P}[\ell(f_{n}(X^{n}))\geq\lfloor\log M_{X}(\beta_{n})\rfloor+\xi]\leq P^{n}(E_{n}^{c})\leq 2^{-n\delta}.$
Hence, by definition, $nR_{n}^{*}(2^{-n\delta},P)\leq\log M_{X}(\beta_{n})+\xi$, and since $\xi\in(0,1)$ was arbitrary,
$R_{n}^{*}(2^{-n\delta},P)\leq\frac{\log M_{X}(\beta_{n})}{n}.$
The remainder of the proof is devoted to showing that there is an appropriate $\beta_{n}$ such that (13) holds, and evaluating the rate $(1/n)\log M_{X}(\beta_{n})$ for this choice.
Let $\alpha\in(0,1)$ arbitrary. We begin by observing that we can express
$M_{X}(\beta_{n})$ $=\sum_{x^{n}\in E_{n}}\frac{P^{n}_{\alpha}(x^{n})}{P^{n}_{\alpha}(x^{n})}$
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp(-\log_{e}P_{\alpha}^{n}(x^{n}))\mathbb{I}_{E_{n}}(x^{n})$
$=\mathbb{E}_{P_{\alpha}^{n}}\Big[\exp\Big(-\sum_{i=1}^{n}\log_{e}P_{\alpha}(X_{i})\Big)\mathbb{I}\Big\{\sum_{i=1}^{n}\log_{e}P(X_{i})\geq\log_{e}\beta_{n}\Big\}\Big].$
We can now apply the version of the Berry-Esséen bound in Lemma 2.2, with $Z_{i}=\log_{e}P_{\alpha}(X_{i})$ and $x=\alpha\log_{e}\beta_{n}-n\log_{e}Z_{\alpha}$, to obtain that:
$M_{X}(\beta_{n})$ $=\mathbb{E}_{P_{\alpha}^{n}}\Big[\exp\Big(-\sum_{i=1}^{n}\log_{e}P_{\alpha}(X_{i})\Big)\mathbb{I}\Big\{\sum_{i=1}^{n}\log_{e}P_{\alpha}(X_{i})\geq\alpha\log_{e}\beta_{n}-n\log_{e}Z_{\alpha}\Big\}\Big]$
$\leq\frac{1}{\sigma_{1,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big)\frac{1}{\sqrt{n}}\exp(-(\alpha\log_{e}\beta_{n}-n\log_{e}Z_{\alpha}))$
$=\frac{1}{\sigma_{1,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big)\frac{Z_{\alpha}^{n}}{\beta_{n}^{\alpha}\sqrt{n}},$ (14)
where $\sigma_{1,\alpha}^{2}=\text{Var}_{P_{\alpha}}(\log_{e}P_{\alpha}(X))$ and $\rho_{1,\alpha}=\mathbb{E}_{P_{\alpha}}|\log_{e}P_{\alpha}(X)-\mathbb{E}_{P_{\alpha}}(\log_{e}P_{\alpha}(X))|^{3}$.
In order to select appropriate values for $\beta_{n}$ and $\alpha$, we examine $P^{n}(E_{n}^{c})$:
$P^{n}(E_{n}^{c})$ $=\sum_{x^{n}\in A^{n}}P^{n}(x^{n})\mathbb{I}_{E_{n}^{c}}(x^{n})$
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(\log_{e}\frac{P^{n}(x^{n})}{P_{\alpha}^{n}(x^{n})}\Bigg)\mathbb{I}\{\log_{e}P^{n}(x^{n})<\log_{e}\beta_{n}\}$
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg)\mathbb{I}\Bigg\{\sum_{i=1}^{n}\log_{e}P(x_{i})<\log_{e}\beta_{n}\Bigg\}.$
Rearranging, we can further express
$P^{n}(E_{n}^{c})$ $=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg)$
$\hskip 72.26999pt\mathbb{I}\Bigg\{n\log_{e}Z_{\alpha}+\sum_{i=1}^{n}\log_{e}P^{1-\alpha}(x_{i})<n\log_{e}Z_{\alpha}+(1-\alpha)\log_{e}\beta_{n}\Bigg\}$
$=\sum_{x^{n}\in A^{n}}P_{\alpha}^{n}(x^{n})\exp\Bigg(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg)$
$\hskip 72.26999pt\mathbb{I}\Bigg\{\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}>-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}\Bigg\},$
so that
$P^{n}(E_{n}^{c})$ $=\mathbb{E}_{P_{\alpha}^{n}}\Bigg\{\exp\Bigg(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(X_{i})}{P(X_{i})}\Bigg)$
$\hskip 72.26999pt\mathbb{I}\Bigg\{\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(X_{i})}{P(X_{i})}>-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}\Bigg\}\Bigg\}.$
Applying Lemma 2.2 again, with $Z_{i}=\log_{e}[P_{\alpha}(X_{i})/P(X_{i})]$ and $x=-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}$, yields
$P^{n}(E_{n}^{c})$ $\leq\frac{1}{\sigma_{2,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big)\frac{1}{\sqrt{n}}\exp\Big(n\log_{e}Z_{\alpha}+(1-\alpha)\log_{e}\beta_{n}\Big)$
$=\frac{1}{\sigma_{2,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big)\frac{Z_{\alpha}^{n}\beta_{n}^{1-\alpha}}{\sqrt{n}}.$ (15)
Now, setting
$\log\beta_{n}=\frac{1}{1-\alpha}\Bigg(-n\delta+\log\Bigg(\frac{1}{\frac{1}{\sigma_{2,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big)\frac{Z_{\alpha}^{n}}{\sqrt{n}}}\Bigg)\Bigg),$
immediately yields (13). Note that, $\beta_{n}>0$ by (15), while (12) and (13) imply $\beta_{n}<1$.
As a result, the upper bound for $M_{X}(\beta_{n})$ in (14) becomes:
$\log M_{X}(\beta_{n})$ $\leq\log\Bigg(\frac{1}{\sigma_{1,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big)\frac{Z_{\alpha}^{n}}{\beta_{n}^{\alpha}\sqrt{n}}\Bigg)$
$=\log\Bigg(\frac{1}{\sigma_{1,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big)\Bigg)+n\log Z_{\alpha}-\frac{1}{2}\log n$
$\hskip 36.135pt+\frac{\alpha n\delta}{1-\alpha}-\frac{\alpha}{1-\alpha}\log\Bigg(\frac{1}{\frac{1}{\sigma_{2,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big)\frac{Z_{\alpha}^{n}}{\sqrt{n}}}\Bigg)$
$=\frac{n}{1-\alpha}(\alpha\delta+\log Z_{\alpha})-\frac{1}{2(1-\alpha)}\log n$
$\hskip 36.135pt+\log\Bigg(\frac{1}{\sigma_{1,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big)\Bigg)+\frac{\alpha}{1-\alpha}\log\Bigg(\frac{1}{\sigma_{2,\alpha}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big)\Bigg).$ (16)
We finally need to select $\alpha$. Since $D(P_{\alpha}\|P)$ is continuous and strictly decreasing in $\alpha$ (Lemma 2.8), and since $\delta\in(D(P_{1}\|P)),D(P_{0}\|P))=(0,D(U\|P))$, there exists a unique $\alpha^{*}\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$. For this value of $\alpha$,
$\alpha^{*}\delta+\log Z_{\alpha^{*}}$ $=\alpha^{*}\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P_{\alpha^{*}}(x)}{P(x)}+\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P_{\alpha^{*}}(x)Z_{\alpha^{*}}}{P_{\alpha^{*}}(x)}$
$=\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P_{\alpha^{*}}(x)^{\alpha^{*}}}{P(x)^{\alpha^{*}}}+\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P(x)^{\alpha^{*}}}{P_{\alpha^{*}}(x)}$
$=\sum_{x\in A}P_{\alpha^{*}}(x)\log\Big(P_{\alpha^{*}}(x)^{(\alpha^{*}-1)}\Big)$
$=(1-\alpha^{*})H(P_{\alpha^{*}}).$ (17)
Substituting this in (16) yields (11), as required.   $\Box$
3.2 Converse
Theorem 3.2 (Converse)
[…]
$R_{n}^{*}(2^{-n\delta},P)\geq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{C}{n},$ (18)
where $P_{\alpha}$ is as defined in (10), $\alpha^{*}$ is the unique $\alpha\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$, and,
$N_{0}$ $=\max\Bigg\{4.4\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1\Big)^{2},\frac{4(1+q+r)^{2}}{p^{2}},\frac{2(1+q+r)}{p(1-\alpha^{*})},N_{1},N_{2}\Bigg\},$
$C$ $=\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big|\tilde{\sigma}_{3}^{2}-(1-\alpha^{*})\hat{\rho}_{3}\Big|+\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3})+1,$
where $\sigma_{3,\alpha}$ and $\rho_{3,\alpha}$ are as defined in Lemma 2.6, and the remaining constants are given by:
$\sigma_{3}^{2}=\sigma_{3,\alpha^{*}}^{2},\;\;\tilde{\sigma}_{3}^{2}=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\hat{\rho}_{3}=\sup_{\alpha\in(0,1)}\rho_{3,\alpha},$
$p=(\log e)(1-\alpha^{*})\sigma_{3}^{2},\;\;q=\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}),\;\;r=19(\log e)(1-\alpha^{*})\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1\Big)\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}},$
$N_{1}=\min\big\{n\geq 8\;:\;\log n\leq p\sqrt{n}\big\},\;\;N_{2}=\min\big\{n\geq 3\;:\;\log n\leq p(1-\alpha^{*})n\big\}.$
Remark. Again, in view of the remark at the end of Section 1.1, the bound (18) remains valid for prefix free codes exactly as stated.
Proof. Taking $k=nR$, for some $R>0$, letting $\tau=n\delta$, and replacing $X$ with $X^{n}$ (and $f^{*}$ with $f_{n}^{*}$) in the one-shot converse in Theorem 2.3, yields:
$\mathbb{P}\Big[\ell(f_{n}^{*}(X^{n}))\geq n\Big(R-\frac{1}{n}\Big)\Big]>\mathbb{P}[-\log P^{n}(X^{n})\geq n(R+\delta)]-2^{-n\delta}.$
Therefore, if we can find a rate $R$ for which
$\mathbb{P}[\log P^{n}(X^{n})\leq-n(R+\delta)]\geq 2\times 2^{-n\delta},$
this would imply that $\mathbb{P}[\ell(f_{n}^{*}(X^{n}))\geq n(R-1/n)]>2^{-n\delta}$, which would in turn give
$R_{n}^{*}(2^{-n\delta},P)\geq R-\frac{1}{n}.$ (19)
Writing
$S_{n}=\{x^{n}\in A^{n}:\log P^{n}(x^{n})\leq-n(R+\delta)\},$
our objective then is to find an appropriate rate $R$ for which
$P^{n}(S_{n})\geq 2\times 2^{-n\delta}.$ (20)
Much of the remainder of the proof will be devoted to establishing (20).
First we will use Lemma 2.4 to express $P^{n}(S_{n})$ in terms of $P_{\alpha_{n}}^{n}(S_{n}^{c})$ for some $\alpha_{n}\in(0,1)$ that will be chosen later. Indeed, since
$S_{n}$ $=\Bigg\{x^{n}\in A^{n}:\sum_{i=1}^{n}\log P^{1-\alpha_{n}}(x_{i})\leq-(1-\alpha_{n})n(R+\delta)\Bigg\}$
$=\Bigg\{x^{n}\in A^{n}:\log\frac{P^{n}(x^{n})}{P^{n}_{\alpha_{n}}(x^{n})}\leq-(1-\alpha_{n})n(R+\delta)+n\log Z_{\alpha_{n}}\Bigg\},$ (21)
we have that
$\frac{P^{n}(x^{n})}{P_{\alpha_{n}}^{n}(x^{n})}\geq\max_{x^{n}\in S_{n}}\frac{P^{n}(x^{n})}{P_{\alpha_{n}}^{n}(x^{n})}\quad\mbox{for all}\;x^{n}\not\in S_{n},$
so we can apply Lemma 2.4 to obtain
$P^{n}(S_{n})=\min_{S\subset A^{n}:P_{\alpha_{n}}^{n}(S^{c})\leq P_{\alpha_{n}}^{n}(S_{n}^{c})}P^{n}(S).$ (22)
The next step will be an application of the Berry-Esséen bound to obtain an upper bound on $P_{\alpha_{n}}^{n}(S_{n}^{c})$. From (21) we get
$P_{\alpha_{n}}^{n}(S_{n}^{c})=P_{\alpha_{n}}^{n}\Bigg(\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha_{n}}(x_{i})}{P(x_{i})}<\frac{(1-\alpha_{n})n(R+\delta)-n\log Z_{\alpha_{n}}}{\log e}\Bigg).$ (23)
Choosing $R$ as
$R=\frac{1}{1-\alpha_{n}}\Big[\mathbb{E}_{P_{\alpha_{n}}}\Big(\log\frac{P_{\alpha_{n}}(X)}{P(X)}\Big)+\log Z_{\alpha_{n}}\Big]-\delta,$ (24)
we have
$\frac{(1-\alpha_{n})n(R+\delta)-n\log Z_{\alpha_{n}}}{\log e}=n\mathbb{E}_{P_{\alpha_{n}}}\Big(\log_{e}\frac{P_{\alpha_{n}}(X)}{P(X)}\Big),$
and the Berry-Esséen in Theorem 2.1 applied to (23) yields
$P_{\alpha_{n}}^{n}(S_{n}^{c})=P_{\alpha_{n}}^{n}\Bigg(\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha_{n}}(x_{i})}{P(x_{i})}<n\mathbb{E}_{P_{\alpha_{n}}}\Big(\log_{e}\frac{P_{\alpha_{n}}(X)}{P(X)}\Big)\Bigg)\leq\Phi(0)+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}.$
Using Lemma 2.6, the bound becomes
$P_{\alpha_{n}}^{n}(S_{n}^{c})\leq\frac{1}{2}+\frac{\frac{\rho_{2,\alpha_{n}}}{(1-\alpha_{n})^{3}}}{\frac{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}{(1-\alpha_{n})^{3}}}=\frac{1}{2}+\frac{\rho_{3,\alpha_{n}}}{2\sigma_{3,\alpha_{n}}^{3}\sqrt{n}}\leq\frac{1}{2}+\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}\sqrt{n}},$ (25)
where $\sigma_{3,\alpha_{n}}$ and $\rho_{3,\alpha_{n}}$ are as in Lemma 2.6, $\tilde{\sigma}_{3}^{2}:=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$ and $\hat{\rho}_{3}:=\sup_{\alpha\in(0,1)}\rho_{3,\alpha}$.
Using this bound, from (22) we get
$P^{n}(S_{n})\geq\min_{S\subset A^{n}:P_{\alpha_{n}}^{n}(S^{c})\leq\frac{1}{2}+\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}\sqrt{n}}}P^{n}(S).$
Then, using the hypothesis testing converse in Theorem 2.5, with $P_{\alpha_{n}}$ in the place of $Q$, $\epsilon=\frac{1}{2}+\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}\sqrt{n}}$, and $\Delta=1$, yields
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big(\frac{1}{2}+\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big)-\frac{1}{2}\log n$ (26)
$\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big(\frac{1}{2}+\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big)-\frac{1}{2}\log n,$ (27)
as long as $\frac{1}{2}+\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}<1$, i.e., for $n>4(\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}+1)^{2}$. Note that $\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}}\geq\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}$ by (25), which justifies (27) and also means that (26) still holds when $n>4(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1)^{2}$.
Recall that the existence of a unique $\alpha^{*}$ as in the statement of the theorem is already established in Theorem 3.1. In order to bound $\log P^{n}(S_{n})$ below further, we will use simple Taylor expansions to get bounds on $D(P_{\alpha_{n}}\|P)$, $\sigma_{2,\alpha_{n}}$ and the $\Phi^{-1}$ term in (27). First we expand $D(P_{\alpha_{n}}\|P)$ around $\alpha^{*}$. Using the expressions for the derivatives of $D(P_{\alpha}\|P)$ in Lemma 2.7, gives
$D(P_{\alpha_{n}}\|P)$ $=D(P_{\alpha^{*}}\|P)+(\alpha_{n}-\alpha^{*})(\alpha^{*}-1)\sigma_{3}^{2}(\log e)$
$\quad+\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Bigg(\sigma_{3,u}^{2}+(u-1)\mathbb{E}_{P_{u}}\Big[\Big(\log_{e}P(X)-\mathbb{E}_{P_{u}}(\log_{e}P(X))\Big)^{3}\Big]\Bigg),$ (28)
for some $u\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha^{*}\leq\alpha_{n}$, and where $\sigma_{3}=\sigma_{3,\alpha^{*}}$. Letting $\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$, we have $\sigma_{3,u}^{2}\leq\hat{\sigma}_{3}^{2}$, and we observe that $|u-1|\leq 1$ and that
$\Big|\mathbb{E}_{P_{u}}\Big[\Big(\log_{e}P(X)-\mathbb{E}_{P_{u}}(\log_{e}P(X))\Big)^{3}\Big]\Big|\leq\mathbb{E}_{P_{u}}\Big|\log_{e}P(X)-\mathbb{E}_{P_{u}}(\log_{e}P(X))\Big|^{3}\leq\hat{\rho}_{3}.$
Therefore,
$D(P_{\alpha_{n}}\|P)\leq D(P_{\alpha^{*}}\|P)+(\alpha_{n}-\alpha^{*})(\alpha^{*}-1)\sigma_{3}^{2}(\log e)+\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}).$ (29)
Next, since $\sigma_{2,\alpha_{n}}=(1-\alpha_{n})\sigma_{3,\alpha_{n}}$ by Lemma 2.6, we can expand $\sigma_{3,\alpha_{n}}$ instead of $\sigma_{2,\alpha_{n}}$ around $\alpha^{*}$, yielding
$\sigma_{3,\alpha_{n}}^{2}=\sigma_{3}^{2}+(\alpha_{n}-\alpha^{*})\mathbb{E}_{P_{w}}\Big[\Big(\log_{e}P(X)-\mathbb{E}_{P_{w}}(\log_{e}P(X))\Big)^{3}\Big]\leq\sigma_{3}^{2}+\hat{\rho}_{3},$
for some $w\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha_{n}-\alpha^{*}\leq 1$. Hence,
$\sigma_{2,\alpha_{n}}=(1-\alpha_{n})\sigma_{3,\alpha_{n}}\leq(1-\alpha^{*})\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}}.$ (30)
And for the third term, we expand $\Phi^{-1}$ around $1/2$ to obtain that
$\Phi^{-1}\Big(\frac{1}{2}+\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big)$ $=\Phi^{-1}\Big(\frac{1}{2}\Big)+\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big)\frac{1}{\phi(\Phi^{-1}(x))}$
$\leq\frac{1}{\sqrt{n}}\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1\Big)\frac{1}{\phi(\Phi^{-1}(\frac{1}{2}+\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}))}$
$\leq\frac{1}{\sqrt{n}}\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1\Big)\frac{1}{\phi(2)}\leq\frac{19}{\sqrt{n}}\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1\Big),$ (31)
for some $x\in(\frac{1}{2},\frac{1}{2}+\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}})$, where (31) holds as long as:
$n\geq 4.4\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1\Big)^{2}>\frac{(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1)^{2}}{(\Phi(2)-1/2)^{2}}.$
Finally, substituting the bounds (29), (30) and (31) into (27), gives
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha^{*}}\|P)-n(\alpha_{n}-\alpha^{*})(\log e)(\alpha^{*}-1)\sigma_{3}^{2}-\frac{n\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3})$
$\hskip 43.36243pt-(1-\alpha^{*})\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}}\times(\log e)\sqrt{n}\times\frac{19}{\sqrt{n}}\Big(\frac{\hat{\rho}_{3}}{\tilde{\sigma}_{3}^{3}}+1\Big)-\frac{1}{2}\log n$
$\geq-n\delta+pn(\alpha_{n}-\alpha^{*})-qn(\alpha_{n}-\alpha^{*})^{2}-r-\frac{1}{2}\log n,$ (32)
where $\delta,p,q$ and $r$ are as in the statement of the theorem.
Now we are in a position to choose $\alpha_{n}$. Letting
$\alpha_{n}=\alpha^{*}+\frac{1}{2p}\frac{\log n}{n}+\frac{1+q+r}{p}\frac{1}{n},$ (33)
simple algebra shows that we have
$pn(\alpha_{n}-\alpha^{*})-qn(\alpha_{n}-\alpha^{*})^{2}-r-\frac{1}{2}\log n\geq 1,$
as long as $n$ is large enough so that
$\frac{\log n}{\sqrt{n}}\leq p\quad\mbox{and}\quad n\geq\frac{4(1+q+r)^{2}}{p^{2}},$ (34)
in which case (32) becomes $\log P^{n}(S_{n})\geq-n\delta+1$, which is exactly our desired bound (20). To ensure its validity, we must also ensure that the earlier conditions we applied – namely, that $\alpha_{n}\geq\alpha^{*}$, that $\alpha_{n}-\alpha^{*}\leq 1$ and that $\alpha_{n}\in(0,1)$ – are also satisfied. The inequalities $\alpha_{n}>0$ and $\alpha_{n}\geq\alpha^{*}$ follow from (33), $\alpha_{n}-\alpha^{*}<1$ follows from (33) and (34), and finally, if we further take $n$ large enough that
$\frac{\log n}{n}<p(1-\alpha^{*})\quad\mbox{and}\quad n>\frac{2(1+q+r)}{p(1-\alpha^{*})},$
then $\alpha_{n}<1$ easily follows from (33). To summarize, taking $n>N_{0}$ as in the statement of the theorem ensures the validity of our key bound (20).
The final step of the proof consists of showing that, with $R$ chosen as in (24) and with $\alpha_{n}$ as in (33), the rate $R$ can be bounded below so that, in combination with (19), we will obtain the desired bound (18) in the statement of the theorem.
To that end, starting from the definition of $R$ in (24), we can re-write
$R$ $=\mathbb{E}_{P_{\alpha_{n}}}\Big(\log\frac{P_{\alpha_{n}}(X)}{P(X)}\Big)+\frac{1}{1-\alpha_{n}}\Big[\alpha_{n}\mathbb{E}_{P_{\alpha_{n}}}\Big(\log\frac{P_{\alpha_{n}}(X)}{P(X)}\Big)+\log Z_{\alpha_{n}}\Big]-\delta$
$=D(P_{\alpha_{n}}\|P)+\frac{1}{1-\alpha_{n}}\Big[\alpha_{n}D(P_{\alpha_{n}}\|P)+(1-\alpha_{n})H(P_{\alpha_{n}})-\alpha_{n}D(P_{\alpha_{n}}\|P)\Big]-\delta$
$=D(P_{\alpha_{n}}\|P)+H(P_{\alpha_{n}})-\delta.$ (35)
To bound this below, we will expand $D(P_{\alpha_{n}}\|P)$ and $H(P_{\alpha_{n}})$ around $\alpha^{*}$. For $D(P_{\alpha_{n}}\|P)$, continuing from the Taylor expansion in (28) we have
$D(P_{\alpha_{n}}\|P)$ $\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Big|\tilde{\sigma}_{3}^{2}-(1-\alpha^{*})\hat{\rho}_{3}\Big|$ (36)
$\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2n}\Big|\tilde{\sigma}_{3}^{2}-(1-\alpha^{*})\hat{\rho}_{3}\Big|$ (37)
$=\delta-\frac{\log n}{2n}-\frac{1+q+r}{n}-\frac{\log e}{2n}\Big|\tilde{\sigma}_{3}^{2}-(1-\alpha^{*})\hat{\rho}_{3}\Big|,$ (38)
where (36) holds because the right-hand side is a lower bound to (28) since $\sigma_{3,u}^{2}\geq\tilde{\sigma}_{3}^{2}$, $1-u\leq 1-\alpha^{*}$ and by the definition of $\hat{\rho}_{3}$; (37) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (38) follows by substituting the expression of $\alpha_{n}$ in (33). Similarly expanding $H(P_{\alpha_{n}})$ around $\alpha^{*}$ and using the expressions for the derivatives of $H(P_{\alpha})$ in Lemma 2.9,
$H(P_{\alpha_{n}})$ $=H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e$
$\hskip 36.135pt-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Bigg(\sigma_{3,v}^{2}+v\mathbb{E}_{P_{v}}\Big[\Big(\log_{e}P(X)-\mathbb{E}_{P_{v}}(\log_{e}P(X))\Big)^{3}\Big]\Bigg)$ (39)
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+1\times\hat{\rho}_{3})$ (40)
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2n}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3})$ (41)
$=H(P_{\alpha^{*}})-\frac{\alpha^{*}}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{\alpha^{*}(1+q+r)}{1-\alpha^{*}}\frac{1}{n}-\frac{\log e}{2n}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}),$ (42)
where (39) holds for some $v\in[\alpha^{*},\alpha_{n}]$; (40) holds since $\sigma_{3,v}^{2}\leq\hat{\sigma}_{3}^{2}$, $v\leq 1$ and by the definition of $\hat{\rho}_{3}$; (41) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (42) follows by substituting the expression of $\alpha_{n}$ in (33).
Substituting the bounds (38) and (42) into (35), gives
$R\geq H(P_{\alpha^{*}})$ $-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}$
$-\Bigg\{\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big|\tilde{\sigma}_{3}^{2}-(1-\alpha^{*})\hat{\rho}_{3}\Big|+\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3})\Bigg\}\times\frac{1}{n},$
and combining this with (19) yields (18) and completes the proof.   $\Box$
4 The number of low-entropy types
In this section we prove a fine combinatorial estimate on the number of stings whose empirical entropy is below a given threshold. This will be important in the proof of the universality result we derive in Section 5, and it may also be of independent interest.
Let $A$ be a fixed finite alphabet of size $|A|=m\geq 2$. We write $\cal P$} for the space of p.m.f.s $P$ on $A$ and denote by ${\cal P}_{n}\subset{\cal P}$ the subset of all $n$-types on $A$; recall that $P\in{\cal P}$ is an $n$-type if, for each $a\in A$, $P(a)$ is of the form $k/n$ for an integer $0\leq k\leq n$. For a string $x^{n}\in A^{n}$, we write $\hat{P}_{x^{n}}$ for the type of $x^{n}$, i.e., for the empirical p.m.f. it induces on $A$:
$\hat{P}_{x^{n}}(a)=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{I}}_{\{x_{i}=a\}},\quad a\in A.$
Clearly, $\hat{P}_{x^{n}}$ is always an $n$-type.
Our main result in this section gives an accurate estimate of the number of strings $x^{n}$ whose types $\hat{P}_{x^{n}}$ have entropy no greater than a given threshold. Recall that the notation $a_{n}=\Theta(b_{n})$ for two nonnegative sequences $a_{n}$ and $b_{n}$ means that there are finite nonzero constants $c,c^{\prime}$ such that $cb_{n}\leq a_{n}\leq c^{\prime}b_{n}$ for all $n$.
Theorem 4.1 (Low-entropy types)
Let $Q$ be a non-uniform p.m.f. of full support on $A$, and for each $n\geq 1$ consider the set
$B_{n}=\big\{x^{n}\in A^{n}:H(\hat{P}_{x^{n}})\leq H(Q)\big\}.$ (43)
Then the cardinality $|B_{n}|$ of $B_{n}$ satisfies:
$|B_{n}|=\Theta\Big(n^{\frac{m-3}{2}}2^{nH(Q)}\Big),$ (44)
where the implied constants depend on $Q,m$.
Theorem 4.1 follows immediately from the lower bound in Theorem 4.4 combined with the upper bound in Theorem 4.6, established in Sections 4.2 and 4.3, respectively. These results also provide additional information on how the implied constants in (44) depend on the problem parameters. In fact, the upper bound in Theorem 4.6 is shown to hold uniformly in a neighbourhood of $Q$.
4.1 Preliminaries
For an $n$-type $P\in{\cal P}_{n}$, we write $\mathcal{T}(P)=\{x^{n}\in A^{n}:\hat{P}_{x^{n}}=P\}$ for the type class of $P$, and we denote the support of any $P\in\mathcal{P}$ as $\mathrm{supp}(P)$.
The following useful notation will be used throughout this section. For two expressions $a$ and $b$ that may depend on $n$, $Q$, and possibly on other parameters, we write $a\gtrsim_{Q}b$ to indicate that that there exists a constant $K=K(Q)$ that depends on $Q$ such that $a\geq K(Q)\cdot b$. Similarly, $a\approx b$ means $a\lesssim b$ and $b\lesssim a$. When no subscripts are given in $\gtrsim,\lesssim$ and $\approx$, the implied constants are absolute. The big-$O$ and small-$o$ notation is used as usual, and again we use subscripts to denote the dependence of the implied constants on the problem parameters. For example, for a positive function $g$, $f(n)=O_{Q}(g(n))$ means that there exists $K=K(Q)$ such that $|f(n)|\leq K(Q)g(n)$ for all $n$ large enough.
The derivations of Theorems 4.4 and 4.6 involve geometric considerations regarding certain subsets of $\cal P$}. We will use the following terminology. A convex body is a convex, open, non-empty, and bounded subset of $\mathbb{R}^{d}$. For a convex body $C,$ we denote its volume by $\mathrm{Vol}(C)$ and its surface area by $\mathrm{Sur}(C).$ The Minkowski sum of two subsets $C_{1},C_{2}$ or ${\mathbb{R}}^{d}$ is $C_{1}+C_{2}=\{x+y:x\in C_{1},y\in C_{2}\}$. Writing $B_{2}(\epsilon)$ for the closed Euclidean ball of radius $\epsilon>0$ centred at the origin, the Minkowski-Steiner formula [14] says that the surface area of a convex body can be expressed as
$\mathrm{Sur}(C)=\lim_{\epsilon\downarrow 0}\frac{\mathrm{Vol}(C+B_{2}(\epsilon))-\mathrm{Vol}(C)}{\epsilon}.$
Moreover, the volume $\mathrm{Vol}(C+B_{2}(\epsilon))$ is locally a polynomial of degree $d$ in $\epsilon$.
A full-rank lattice $\Gamma$ in ${\mathbb{R}}^{d}$ is a discrete additive subgroub $\Gamma$ of ${\mathbb{R}}^{d}$ generated by $d$ linearly independent elements $v_{1},\ldots,v_{d}$ of ${\mathbb{R}}^{d}$. The lemma below is a standard result on counting lattice points in a convex body; see [35, Lemma 3.22].
Lemma 4.2
Let $\Gamma\subset\mathbb{R}^{d}$ be a full-rank lattice generated by $\{v_{1},\ldots,v_{d}\}$, and let $B\subset\mathbb{R}^{d}$ be a convex body. Then for $R>0$,
$|(R\cdot B)\cap\Gamma|=(R^{d}+O_{\Gamma,B,d}(R^{d-1}))\frac{\operatorname{Vol}(B)}{\mathrm{Vol}(v_{1}\wedge\ldots\wedge v_{d})},$
where $v_{1}\wedge\ldots\wedge v_{d}$ denotes the parallelepiped with vertices $v_{1},\ldots,v_{d}$, and the $O_{\Gamma,B,d}(R^{d-1})$ term is bounded in absolute value by $KR^{d-1}$, for some constant $K$ that depends on $\Gamma,B$ and $d$.
The next lemma is a standard result, which can be obtained by using the Robbins refinement of the Stirling approximation in the multinomial coefficient; see, e.g., [10].
Lemma 4.3 (Size of type classes)
Let $P$ be an $n$-type with full support on an alphabet of size $k$. Then
$|{\cal T}(P)|\approx_{k}\frac{2^{nH(P)}}{n^{(k-1)/2}}\,\prod_{a\in A}\frac{1}{\sqrt{P(a)}},$
where the implied constants depend only on $k$.
4.2 Lower bound
Theorem 4.4
Let $Q$ be a non-uniform p.m.f. of full support on $A$. Then the set $B_{n}$ in (43) satisfies,
$\big|\big\{x^{n}:H(\hat{P}_{x^{n}})\leq H(Q)\big\}\big|\gtrsim_{Q,m}2^{nH(Q)}n^{\frac{(m-3)}{2}},$
for all $n\geq N(Q,m)$, for an integer $N(Q,m)$ that depends on $Q$ and $m$ only.
A key element of the proof of Theorem 4.4 is the following estimate.
Lemma 4.5
Under the assumptions of Theorem 4.4, we have
$\Big|\Big\{P\in\mathcal{P}_{n}:H(P)\in\Big[H(Q)-\frac{1}{n},H(Q)\Big]\Big\}\Big|\gtrsim_{Q,m}n^{m-2}+n^{m-3},$
for all $n\geq N^{\prime}(Q,m)$, for an integer $N^{\prime}(Q,m)$ that depends on $Q$ and $m$ only.
Proof. Let d=m1d=m-1. We identify each $y$ in the simplex
$\mathcal{S}=\Big\{z\in\mathbb{R}^{d}:\sum_{i=1}^{d}z_{i}\leq 1,z_{i}\geq 0\Big\},$
with a p.m.f. $P_{y}\in\mathcal{P}$ by setting $P_{y}=(y,1-\sum_{i=1}^{d}y_{i})\in\mathbb{R}^{m}$. We define the following sets in $\mathbb{R}^{d}$:
${\cal A}_{Q,n}$ $=\Big\{y\in\mathcal{S}:H(P_{y})>H(Q)-\frac{1}{n}\Big\}$
${\cal A}_{Q}$ $=\{y\in\mathcal{S}:H(P_{y})>H(Q)\}.$
The set of all $n$-types $P\in\mathcal{P}_{n}$ such that $H(P)\in[H(Q)-1/n,H(Q)]$ has the same cardinality (up to constants, since we define ${\cal A}_{Q,n}$ with >> rather than $\geq$ in order for it to be open) as the set of lattice points
$\Big(\frac{1}{n}\cdot\mathbb{Z}^{d}\Big)\cap({\cal A}_{Q,n}\cap{\cal A}_{Q}^{c})$ $=\Big(\frac{1}{n}\cdot\mathbb{Z}^{d}\Big)\cap{\cal A}_{Q,n}-\Big(\frac{1}{n}\cdot\mathbb{Z}^{d}\Big)\cap{\cal A}_{Q}$
$=\mathbb{Z}^{d}\cap\bigl({n}\cdot{\cal A}_{Q,n}\bigr)-\mathbb{Z}^{d}\cap\bigl({n}\cdot{\cal A}_{Q}\bigr).$ (45)
Note that by the concavity of entropy, both ${\cal A}_{Q}$ and ${\cal A}_{Q,n}$ are convex bodies. Our goal is to use Lemma 4.2 with $\Gamma=\mathbb{Z}^{d}$, generated by $\{v_{1},\ldots,v_{d}\}=\{e_{1},\ldots,e_{d}\}$ being the standard basis in ${\mathbb{R}}^{d}$, and with $R=n$, in order to estimate $\big|\mathbb{Z}^{d}\cap\bigl({n}\cdot{\cal A}_{Q,n}\bigr)\big|$ and $\big|\mathbb{Z}^{d}\cap\bigl({n}\cdot{\cal A}_{Q}\bigr)\big|$ separately.
For the latter, Lemma 4.2 applies directly with $B={\cal A}_{Q}$ to give
$\big|\mathbb{Z}^{d}\cap\bigl({n}\cdot{\cal A}_{Q}\bigr)\big|=\Big(n^{d}+O_{Q,d}(n^{d-1})\Big)\mathrm{Vol}({\cal A}_{Q}).$ (46)
Since $Q$ has full support and is not the uniform distribution, the gradient of the entropy functional, $y\mapsto H(P_{y})$, $y\in\mathcal{S}$ is bounded above near the boundary of ${\cal A}_{Q}$, and thus we may apply a multivariate, first-order Taylor expansion to see that, for $n$ large enough depending on $Q$, and for some appropriately small constant $K(Q)$ that depends on $Q$ only,
${\cal A}_{Q}+B_{2}\Bigl(\frac{K(Q)}{n}\Bigr)\subset{\cal A}_{Q,n}.$ (47)
Therefore,
$\big|\mathbb{Z}^{d}\cap\bigl(n\cdot{\cal A}_{Q,n}\bigr)\big|\geq\Big|\mathbb{Z}^{d}\cap\Bigl[n\cdot\Big({\cal A}_{Q}+B_{2}\Big(\frac{K(Q)}{n}\Big)\Big)\Bigr]\Big|.$ (48)
Now we claim that
$\Big|\mathbb{Z}^{d}\cap\Bigl[n\cdot\Big({\cal A}_{Q}+B_{2}\Big(\frac{K(Q)}{n}\Big)\Big)\Bigr]\Big|\geq\Bigl(n^{d}+O_{Q,d}(n^{d-1})\Bigr)\mathrm{Vol}\Bigl({\cal A}_{Q}+B_{2}\Big(\frac{K(Q)}{n}\Big)\Bigr).$ (49)
Lemma 4.2 does not apply directly, since the error term depends on the set $B$ and we wish to take $B={\cal A}_{Q}+B_{2}(K(Q)/n)$, which depends on $n$. Nevertheless, examining the proof in [35], we see that that the error term is – up to constants depending only on the dimension – the volume of the surface of $R\cdot B=n\cdot\bigl({\cal A}_{Q}+B_{2}(K(Q)/n)\bigr)$, which is still of the same order, i.e.,
$\mathrm{Sur}\left(n\cdot\Big({\cal A}_{Q}+B_{2}\Big(\frac{K(Q)}{n}\Big)\Big)\right)=O_{Q,d}{(n^{d-1})}.$ (50)
Thus, repeating the proof of Lemma 4.2 as in [35], but replacing the estimate of the error term with (50), we conclude that (49) holds.
Combining (45) with the estimates (46) and (49), we conclude that
$\Big|\Big\{P\in\mathcal{P}_{n}:H(P)\in\ \Big[H(Q)-\frac{1}{n},H(Q)\Big]\Big\}\Big|$
$\geq\big(n^{d}+O_{Q,d}(n^{d-1})\big)\Bigl[\mathrm{Vol}\Bigl({\cal A}_{Q}+B_{2}\Big(\frac{K(Q)}{n}\Big)\Bigr)-\mathrm{Vol}({\cal A}_{Q})\Bigr]$ (51)
$\gtrsim_{Q,d}n^{d-1}+n^{d-2}$ (52)
for $n$ large enough depending on $Q$ and $d$, where the implied constant in the last inequality depends to first order on $K(Q)$ and the surface area of $A_{Q}$.
The result follows upon replacing $d$ by m1m-1.   $\Box$
Proof of Theorem $4.$ By Lemma 4.3, for any $n$-type $P$ with full support and $H(Q)-\frac{1}{n}\leq H(P)\leq H(Q)$,
$|\mathcal{T}(P)|\gtrsim_{Q,m}n^{-\frac{m-1}{2}}2^{nH(Q)}.$ (53)
Therefore,
$|\{x^{n}:H(\hat{P}_{x^{n}})\leq H(Q)\}|$
$\geq\sum_{P:H(P)\leq H(Q)}|\mathcal{T}(P)|$
$\geq\sum_{P:H(Q)-\frac{1}{n}\leq H(P)\leq H(Q)}|\mathcal{T}(P)|$
$\gtrsim_{Q,m}n^{-\frac{m-1}{2}}2^{nH(Q)}\Big|\Big\{P\in\mathcal{P}_{n}:H(P)\in\Big[H(Q)-\frac{1}{n},H(Q)\Big]\Big\}\Big|,$ (54)
where in the last step we used (53).
Combining (54) with Lemma 4.5 we obtain that, for $n$ large enough depending on $Q$ and $m$,
$|\{x^{n}:H(\hat{P}_{x^{n}})\leq H(Q)\}|\gtrsim_{Q,m}n^{m-2}\frac{1}{n^{\frac{m-1}{2}}}2^{nH(Q)}=n^{\frac{m-3}{2}}2^{nH(Q)},$
which is the claimed result.   $\Box$
4.3 Upper bound
Next we establish a matching upper bound. In fact, we prove a stronger result, where the polynomial prefactor, the implied constants, and the value of $n$ beyond which the bound holds are the same for all $Q^{\prime}$ a neighbourhood of $Q$.
Theorem 4.6
Let $Q$ be a non-uniform p.m.f. of full support on $A$. Suppose $\{Q_{n}\}_{n\geq 1}$ be a sequence of p.m.f.s on $A$, such that $Q_{n}\to Q$ as $n\to\infty$. Then
$\big|\{x^{n}\in A^{n}:\ H(\hat{P}_{x^{n}})\leq H(Q_{n})\}\ \big|\lesssim_{Q,m}\ n^{\frac{m-3}{2}}\,2^{nH(Q_{n})},$
for all $n\geq n_{0}(Q,m)$, where $n_{0}(Q,m)$ depends only on $Q$ and $m$.
Proof. Assume $n$ is large enough such that $Q_{n}$ is not uniform and has full support. We have
$\big|\{x^{n}\in A^{n}:\ H(\hat{P}_{x^{n}})\leq H(Q_{n})\}\big|=\sum_{r=1}^{m}\sum_{j\geq 0}\sum_{P\in{\cal S}_{j,r}}|\mathcal{T}(P)|,$ (55)
where, for $r\in\{1,\dots,m\}$ and $j\in\mathbb{Z}_{+}$,
${\cal S}_{j,r}:=\Bigl\{P\in{\cal P}_{n}:\ |\mathrm{supp}(P)|=r,\ \frac{j}{n}\leq H(Q_{n})-H(P)<\frac{j+1}{n}\Bigr\}.$
Applying Lemma 4.3 and noting that on ${\cal S}_{j,r}$ we have $2^{nH(P)}\leq 2^{nH(Q_{n})}\,2^{-j}$, we see that
$\sum_{P\in{\cal S}_{j,r}}|{\cal T}(P)|\ \lesssim_{r}\frac{2^{nH(Q_{n})}\,2^{-j}}{n^{(r-1)/2}}\ \sum_{P\in{\cal S}_{j,r}}\ \prod_{a\in\mathrm{supp}(P)}\frac{1}{\sqrt{P(a)}}.$ (56)
Setting
$J_{n}:=\Big\lceil(m+3)\,\log_{2}n\Big\rceil,$ (57)
we split the inner sum in (55) as
$\sum_{j\leq J_{n}}\ \sum_{P\in{\cal S}_{j,r}}|{\cal T}(P)|\quad+\quad\sum_{j>J_{n}}\ \sum_{P\in{\cal S}_{j,r}}|{\cal T}(P)|.$ (58)
First we show that second term in (58) has negligible contribution due to the exponential factor. For a fixed $r$ and $j>J_{n}$ we have
$|{\cal S}_{j,r}|\leq\ |\{P\in{\cal P}_{n}:|\mathrm{supp}(P)|=r\}|\lesssim n^{r-1},$
and for any type $P$ with $r$ positive coordinates,
$\prod_{a\in\mathrm{supp}(P)}\tfrac{1}{\sqrt{P(a)}}\leq(\sqrt{n})^{\,r}=n^{r/2},$
since each $P(a)\geq 1/n$ whenever it is positive. Substituting these into (56) yields
$\sum_{P\in{\cal S}_{j,r}}|{\cal T}(P)|\ \lesssim_{m}\,2^{nH(Q_{n})}\,2^{-j}\,n^{\,r-\frac{1}{2}},$
where we noted that the dependence of the implied constant on $r$ is equivalent to dependence on $m$. Hence,
$\sum_{j>J_{n}}\ \sum_{P\in{\cal S}_{j,r}}|{\cal T}(P)|\$$\lesssim_{Q,m}2^{nH(Q_{n})}\,n^{\,r-\frac{1}{2}}\sum_{j>J_{n}}2^{-j}$
$\lesssim_{Q,m}2^{nH(Q_{n})}\,n^{\,r-\frac{1}{2}}\,2^{-J_{n}}$
$=o_{Q,m}\big(2^{nH(Q_{n})}\,n^{\frac{m-3}{2}}\big),$
where we used (57) and the fact that $r\leq m$. Summing over $r$ we conclude that
$\sum_{r=1}^{m}\sum_{j>J_{n}}^{n-1}\sum_{P\in{\cal S}_{j,r}}|\mathcal{T}(P)|=o_{Q,m}\big(2^{nH(Q_{n})}\,n^{\frac{m-3}{2}}\big).$ (59)
We now bound the first term in (58), corresponding to $j\leq J_{n}$. Let $\delta=\delta(Q)>0$ to be chosen later. For notational convenience, for the rest of the proof we take, without loss of generality, the alphabet $A$ to be $\{1,2,\ldots,m\}$. For $1\leq r\leq m$ and $s=0,1,\dots,r$, let
${\cal S}_{j,r}^{(s)}:=\Big\{P\in{\cal S}_{j,r}:|\{i:\ P(i)\leq\delta\}|=s\Big\},$
and write
$F(P):=\{i:\ P(i)\leq\delta\},\quad\mbox{so that}\;|F(P)|=s\;\mbox{for}\;P\in S_{j,r}^{(s)}.$
We bound $\sum_{P\in{\cal S}_{j,r}}|{\cal T}(P)|$ by splitting over $s$ and $F$. Clearly, since $j\leq J_{n}$ and $J_{n}/n\to 0$, by choosing $\delta$ small enough depending only on $Q$ and taking $n$ large enough so that $J_{n}/n$ is small and $H(Q_{n})$ is close to $H(Q)$, there have to be at least two coordinates of any $P\in{\cal S}_{j,r}^{(s)}$ with $P(i)>\delta$ (as, otherwise, $H(P)$ would be smaller than $H(Q_{n})-\epsilon$ for some $\epsilon>0$). Therefore, it suffices to restrict to $s\leq r-2$:
$\sum_{P\in{\cal S}_{j,r}}|{\cal T}(P)|=\sum_{s=0}^{r-2}\ \sum_{\begin{subarray}{c}F\subseteq A\\ |F|=s\end{subarray}}\ \sum_{P\in{\cal S}_{j,r}^{(s)}:\ F(P)=F}|{\cal T}(P)|.$
By Lemma 4.3 and the fact that $2^{nH(P)}\leq 2^{nH(Q_{n})}\,2^{-j}$ on ${\cal S}_{j,r}$,
$|{\cal T}(P)|\ \lesssim_{m}\frac{2^{nH(Q_{n})}\,2^{-j}}{n^{(r-1)/2}}\ \prod_{i=1}^{r}\frac{1}{\sqrt{P(i)}}.$
Hence,
$\sum_{P\in{\cal S}_{j,r}^{(s)}:\ F(P)=F}|{\cal T}(P)|$ $\lesssim_{m}\frac{2^{nH(Q_{n})}\,2^{-j}}{n^{(r-1)/2}}\sum_{P\in{\cal S}_{j,r}^{(s)}:\ F(P)=F}\Big(\prod_{i\notin F}\frac{1}{\sqrt{P(i)}}\Big)\Big(\prod_{i\in F}\frac{1}{\sqrt{P(i)}}\Big)$
$\lesssim_{m,\delta}\frac{2^{nH(Q_{n})}\,2^{-j}}{n^{(r-1)/2}}\sum_{P\in{\cal S}_{j,r}^{(s)}:\ F(P)=F}\ \prod_{i\in F}\frac{1}{\sqrt{P(i)}},$ (60)
where we used $P(i)\geq\delta$ for $i\notin F$, so $\prod_{i\notin F}P(i)^{-1/2}\leq\delta^{-(r-s)/2}$.
Now parametrise the $s$ “small” coordinates by their counts $c_{i}:=nP(i)\in\{1,2,\dots,\lfloor\delta n\rfloor\}$ (note that, as before, $c_{i}\geq 1$ whenever $P(i)>0$). After fixing these counts, the remaining rsr-s coordinates then live on a compact subset of $\mathbb{R}^{r-s-1}$. Moreover, the entropy of the type has to be concentrated on those coordinates, as the entropy contribution of the rest can be made arbitrarily small by choosing $\delta(Q)$ small enough. Thus, we can use an analogous argument to (47)–(52), but with the corresponding upper bounds, to conclude that for $n$ large enough depending on $Q$ and $m$ there are
[…]
Note that ${\cal A}_{Q_{n}}$ is still a convex body and its boundary is a smooth hypersurface, so we still have
$\mathrm{Sur}\left(n\cdot\Big({\cal A}_{Q_{n}}+B_{2}\Big(\frac{K(Q)}{n}\Big)\Big)\right)=O_{Q,d}{(n^{d-1})}.$
[In fact, the implied constant in this estimate can be taken to be independent of $Q$, since ${\cal A}_{Q_{n}}$ is a convex subset of the simplex]. The reverse inclusion to (47) can be justified by observing that the entropy is strictly concave, and since $Q$ is not the uniform distribution, the gradient is also uniformly bounded below coordinatewise near the boundary of ${\cal A}_{Q_{n}}$. The analogous upper bound to (52) can also be justified in the same way as (52), since $\mathrm{Vol}(C+B_{2}(\epsilon))$ is locally polynomial in $\epsilon$ for any convex body $C$.
Therefore, by (61), for $n$ large enough depending on $Q$ and $m,$
$\sum_{P\in{\cal S}_{j,r}^{(s)}:\ F(P)=F}\ \prod_{i\in F}\frac{1}{\sqrt{P(i)}}$ $\lesssim_{Q,m}\sum_{(c_{i})_{i\in F}}\ \Big(\prod_{i\in F}\sqrt{\frac{n}{c_{i}}}\Big)\ \cdot n^{\,r-s-2}$
$\leq\Big(\prod_{i\in F}\sum_{c=1}^{\lfloor\delta n\rfloor}\sqrt{\frac{n}{c}}\Big)\ \cdot n^{\,r-s-2}$
$\lesssim_{\delta,m}n^{\,s}\,n^{\,m-s-2},$ (62)
since we obviously have
$\sum_{c=1}^{\lfloor\delta n\rfloor}\sqrt{n/c}\leq\sqrt{n}\sum_{c=1}^{\lfloor\delta n\rfloor}c^{-1/2}\lesssim_{\delta}n.$
Combining (60) and (62), and summing over all $\binom{r}{s}$ choices for $F$, we get, for $n$ large enough depending on $Q,m,$
$\sum_{P\in{\cal S}_{j,r}^{(s)}}|{\cal T}(P)|$ $\lesssim_{Q,m}2^{nH(Q_{n})}\,2^{-j}{n^{-(r-1)/2}}\ \cdot\ n^{\,r-s-2}\,n^{\,s}$
$=2^{nH(Q_{n})}\,2^{-j}\,n^{\frac{r-3}{2}}$
$\leq 2^{nH(Q_{n})}\,2^{-j}\,n^{\frac{m-3}{2}},$
since $r\leq m$. Summing over $1\leq r\leq m$, $0\leq s\leq m-2$, and $j\leq J_{n}$ yields:
$\sum_{r=1}^{m}\sum_{j\leq J_{n}}\ \sum_{P\in{\cal S}_{j,m}}|{\cal T}(P)|\lesssim_{Q,m}2^{nH(Q_{n})}\,n^{\frac{m-3}{2}}.$ (63)
Finally, combining (59) and (63) yields the claimed bound.   $\Box$
5 Universal compression
Theorem 5.1 (Universal achievability)
[…]
${\mathbb{P}}(\ell(\phi_{n}^{*}(X^{n}))\geq nR^{\sf u}(n,\delta,P)\big)\leq 2^{-n\delta}\quad\mbox{eventually},$ (64)
at a rate $R^{\sf u}(n,\delta,P)$ that satisfies, as $n\to\infty$,
$nR^{\sf u}(n,\delta,P)\leq nH(P_{\alpha^{*}})+\Big(\frac{m-2}{2}-\frac{1}{2(1-\alpha^{*})}\Big)\log n+O(1),$ (65)
where $\alpha^{*}$ is the unique $\alpha\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$, the p.m.f. $P_{\alpha}$ is defined in (10), and the implied constant in the $O(1)$ term depends on $m,P$ and $\delta$.
Remark. An examination of the proof below shows that the excess-rate probability bound in (64) holds for all $n$ greater than some $M_{0}(P,\delta)$, which can be explicitly identified in terms of $P$ and $\delta$ in a similar way to how the expression for $N_{0}$ was derived in Theorem 3.2. But in the present context, the exact value of $M_{0}(P,\delta)$ is less relevant, since our bound on the rate in (65) is already an asymptotic statement.
The observation in the following lemma, which follows directly from the definitions, will be useful in the proof of Theorem 5.1.
Lemma 5.2
For all $\alpha\in(0,1)$, and any pair of p.m.f.s $P,Q$ on $A$ such that $P$ has full support, we have:
$\alpha[D(Q\|P)-D(P_{\alpha}\|P)]=D(Q\|P_{\alpha})+(1-\alpha)[H(Q)-H(P_{\alpha})].$
Proof of Theorem 5.1. The construction of the compressors $\{\phi_{n}^{*}\}$ is very simple. For each $n$, all strings $x^{n}\in A^{n}$ are sorted in order of increasing empirical entropy $H(\hat{P}_{x^{n}})$, with ties broken arbitrarily. This ordering is shared between the encoder and decoder. Then, each $x^{n}$ is described by the binary representation of its index in the ordered list, so that the $k$th string $x^{n}$ has a description length of $\ell(\phi_{n}^{*}(x^{n}))=\lfloor\log k\rfloor$ bits.
Choose and fix a p.m.f. with full support on $A$ and a $\delta\in(0,D(U\|P))$. As in Theorems 3.1 and 3.2, let $\alpha^{*}$ be the unique $\alpha\in(0,1)$ such that $D(P_{\alpha^{*}}\|P)=\delta$. For each $n\geq 1$, let $\alpha_{n}\in(0,1)$ be a constant to be chosen later. Consider the sets
$E_{n}=\{x^{n}\in A^{n}:H(\hat{P}_{x^{n}})\leq H(P_{\alpha_{n}})\},$
and let $R^{\sf u}(n,\delta,P)=\frac{1}{n}[\log|E_{n}|+1]$.
We first bound the excess-rate probability of $\phi_{n}^{*}$. This step will take up most of the proof. Specifically, we show that there is a sequence $\{\alpha_{n}\}$ such that, for all $n$ large enough, $\alpha_{n}\in[\alpha^{*},1)$ and
${\mathbb{P}}\big(\ell(\phi_{n}^{*}(X^{n}))\geq nR^{\sf u}(n,\delta,P)\big)={\mathbb{P}}\big(\ell(\phi_{n}^{*}(X^{n}))\geq\log|E_{n}|+1\big)\leq P^{n}(E_{n}^{c})\leq 2^{-n\delta},$
which implies (64). The equality and the first inequality follow from the definitions, so we need to show that
$P^{n}(E_{n}^{c})\leq 2^{-n\delta}\quad\mbox{eventually}.$ (66)
To that end, we begin by expressing
$P^{n}(E_{n}^{c})=\sum_{x^{n}\in A^{n}}P^{n}(x^{n})\mathbb{I}_{E_{n}^{c}}(x^{n})=\sum_{x^{n}\in A^{n}}P_{\alpha^{*}}^{n}(x^{n})\exp\Big(\log_{e}\frac{P^{n}(x^{n})}{P_{\alpha^{*}}^{n}(x^{n})}\Big)\mathbb{I}\{H(\hat{P}_{x^{n}})>H(P_{\alpha_{n}})\}.$
Noting that, for every $x^{n}$,
$H(\hat{P}_{x^{n}})=\frac{1}{n}\sum_{i=1}^{n}\log\frac{1}{P(x_{i})}-D(\hat{P}_{x^{n}}\|P),$
we can further express $P^{n}(E_{n}^{c})$ as
$P^{n}(E_{n}^{c})$ $=\sum_{x^{n}\in A^{n}}P_{\alpha^{*}}^{n}(x^{n})\exp\Big(\log_{e}\frac{P^{n}(x^{n})}{P_{\alpha^{*}}^{n}(x^{n})}\Big)$
$\hskip 57.81621pt\mathbb{I}\left\{H(\hat{P}_{x^{n}})>H(P_{\alpha_{n}}),\;\frac{1}{n}\sum_{i=1}^{n}\log\frac{1}{P(x_{i})}>H(P_{\alpha_{n}})+D(\hat{P}_{x^{n}}\|P)\right\}$
$\leq\sum_{x^{n}\in A^{n}}P_{\alpha^{*}}^{n}(x^{n})\exp\Big(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha^{*}}(x_{i})}{P(x_{i})}\Big)\mathbb{I}\Bigg\{\frac{1}{n}\sum_{i=1}^{n}\log\frac{1}{P(x_{i})}>H(P_{\alpha_{n}})+D(P_{\alpha_{n}}\|P)\Bigg\},$
where the last inequality follows from Lemma 5.2. Recalling the definition of $Z_{\alpha}$ from (10) and defining $\beta_{n}>0$ via $H(P_{\alpha_{n}})+D(P_{\alpha_{n}}\|P)=-\frac{\log\beta_{n}}{n}$, we have
$P^{n}(E_{n}^{c})$ $\leq\sum_{x^{n}\in A^{n}}\Bigg[P_{\alpha^{*}}^{n}(x^{n})\exp\Big(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha^{*}}(x_{i})}{P(x_{i})}\Big)$
$\hskip 72.26999pt\mathbb{I}\Bigg\{n\log_{e}Z_{\alpha^{*}}+\sum_{i=1}^{n}\log_{e}P^{1-\alpha^{*}}(x_{i})<n\log_{e}Z_{\alpha^{*}}+(1-\alpha^{*})\log_{e}\beta_{n}\Bigg\}\Bigg]$
$=\mathbb{E}_{P_{\alpha^{*}}^{n}}\Bigg[\exp\Big(-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha^{*}}(X_{i})}{P(X_{i})}\Big)$
$\hskip 72.26999pt\mathbb{I}\Bigg\{\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha^{*}}(X_{i})}{P(X_{i})}>-n\log_{e}Z_{\alpha^{*}}-(1-\alpha^{*})\log_{e}\beta_{n}\Bigg\}\Bigg].$
We further bound this expression using the Berry-Esséen bound in the form given in Lemma 2.2, with $Z_{i}=\log_{e}[P_{\alpha^{*}}(X_{i})/P(X_{i})]$ and $x=-n\log_{e}Z_{\alpha^{*}}-(1-\alpha^{*})\log_{e}\beta_{n}$, to obtain that
$P^{n}(E_{n}^{c})$ $\leq\frac{1}{\sigma_{2}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2}}{\sigma_{2}^{2}}\Big)\frac{1}{\sqrt{n}}\exp\Big(n\log_{e}Z_{\alpha^{*}}+(1-\alpha^{*})\log_{e}\beta_{n}\Big)$
$=\frac{1}{\sigma_{2}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2}}{\sigma_{2}^{2}}\Big)\frac{Z_{\alpha^{*}}^{n}\beta_{n}^{1-\alpha^{*}}}{\sqrt{n}},$ (67)
where $\sigma_{2}$ and $\rho_{2}$ are as in Theorem 3.1. Let $\bar{r}=\frac{1}{1-\alpha^{*}}\log\big(\frac{1}{\sigma_{2}}(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2}}{\sigma_{2}^{2}})\big)$. If $\beta_{n}$ satisfies
$\log\beta_{n}$ $\leq\frac{1}{1-\alpha^{*}}\Bigg[-n\delta-\log\Bigg(\frac{1}{\sigma_{2}}\Big(\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2}}{\sigma_{2}^{2}}\Big)\frac{Z_{\alpha^{*}}^{n}}{\sqrt{n}}\Bigg)\Bigg]$
$=-\frac{1}{1-\alpha^{*}}(\delta+\log Z_{\alpha^{*}})n+\frac{1}{2(1-\alpha^{*})}\log n-\bar{r},$ (68)
then (67) immediately implies that (66) holds, which gives (64), as required.
So it only remains to show that $\{\alpha_{n}\}$ can be selected so that (68) holds. In order to choose $\{\alpha_{n}\}$ appropriately, we recall that $\alpha_{n}\geq\alpha^{*}$ eventually and that the Taylor expansion (36) in the proof of Theorem 3.2 gives, for all $n$ large enough,
$D(P_{\alpha_{n}}\|P)\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Big|\tilde{\sigma}_{3}^{2}-(1-\alpha^{*})\hat{\rho}_{3}\Big|,$ (69)
where the constants $\sigma_{3}$, $\tilde{\sigma}_{3}$ and $\hat{\rho}_{3}$ are defined in Lemma 2.6 and Theorem 3.2. Similarly, the Taylor expansion (40) in the proof of Theorem 3.2 gives, for all $n$ large enough,
$H(P_{\alpha_{n}})\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}),$ (70)
where the constant $\hat{\sigma}_{3}$ is defined in Theorem 3.2. Therefore, combining (69) with (70), and recalling that $\delta=D(P_{\alpha^{*}}\|P)$, we have
$H(P_{\alpha_{n}})+D(P_{\alpha_{n}}\|P)$ $\geq H(P_{\alpha^{*}})+\delta-(\alpha_{n}-\alpha^{*})\bar{p}-(\alpha_{n}-\alpha^{*})^{2}\bar{q},$ (71)
where $\bar{p}=\sigma_{3}^{2}(\log e)>0$ and $\bar{q}=\frac{\log e}{2}\big(|\tilde{\sigma}_{3}^{2}-(1-\alpha^{*})\hat{\rho}_{3}|+\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}\big)>0$.
Now we are in a position to choose $\{\alpha_{n}\}$. Letting, for each $n$,
$\alpha_{n}=\alpha^{*}+\frac{1}{2\bar{p}(1-\alpha^{*})}\frac{\log n}{n}-\frac{\bar{q}+\bar{r}}{\bar{p}}\frac{1}{n},$
we see that $1>\alpha_{n}\geq\alpha^{*}$ eventually. Also, simple algebra shows that, eventually,
$(\alpha_{n}-\alpha^{*})\bar{p}+(\alpha_{n}-\alpha^{*})^{2}\bar{q}\leq\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{\bar{r}}{n},$
which, recalling from (17) that $H(P_{\alpha^{*}})=\frac{1}{1-\alpha^{*}}(\alpha^{*}\delta+\log Z_{\alpha^{*}})$, gives,
$H(P_{\alpha^{*}})+\delta-(\alpha_{n}-\alpha^{*})\bar{p}-(\alpha_{n}-\alpha^{*})^{2}\bar{q}\geq\frac{1}{1-\alpha^{*}}(\delta+\log Z_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}+\frac{\bar{r}}{n}.$ (72)
Combining (71) with (72) and the definition of $\beta_{n}$ implies that (68) holds, which gives the required bound (64) on the excess-rate probability.
Finally, we will establish the claimed bound (65) on the rate, $R^{\sf u}(n,\delta,P)=\frac{1}{n}[\log|E_{n}|+1]$. The key step in the evaluation of $R^{\sf u}(n,\delta,P)$ was already established in Theorem 4.6, from which we obtain that, as $n\to\infty$,
$\log|E_{n}|\leq nH(P_{\alpha_{n}})+\Big(\frac{m-3}{2}\Big)\log n+O(1).$ (73)
Also, from the earlier Taylor expansion in (39) we can bound, for $n$ large enough,
$H(P_{\alpha_{n}})$ $\leq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e+\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3})$ (74)
$\leq H(P_{\alpha^{*}})-\Big(\frac{1}{2\bar{p}(1-\alpha^{*})}\frac{\log n}{n}-\frac{\bar{q}+\bar{r}}{n}\Big)\alpha^{*}\bar{p}+\frac{\log e}{2n}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}),$
where (74) holds since $\sigma_{3,v}^{2}\leq\hat{\sigma}_{3}^{2}$, $v\leq 1$, and by the definition of $\hat{\rho}_{3}$. Therefore, as $n\to\infty$,
$H(P_{\alpha_{n}})\leq H(P_{\alpha^{*}})-\frac{\alpha^{*}}{2(1-\alpha^{*})}\frac{\log n}{n}+O\Big(\frac{1}{n}\Big),$ (75)
and combining (73) with (75) yields the required bound (65) on the rate and completes the proof.   $\Box$
6 Concluding remarks
This work revisits the problem of lossless data compression, in the case when strict guarantees are required on the probability that the target compression rate might be exceeded. When this excess-rate probability is required to be very small, classical approximations to the best achievable target rate – à la Strassen – fail to be informative. Instead, much more appropriate approximations are obtained by examining the best achievable rate when the excess-rate probability is exponentially small in the blocklength. The new approximations to the optimal rate are derived from explicit nonasymptotic bounds, that are accurate up to and including terms of order $\frac{\log n}{n}$. These bounds support a crucial operational conclusion: In applications where the blocklength is only moderately large and where stringent guarantees are required on the rate, the best achievable rate is no longer close to the entropy. Instead, it is given by the more pragmatic rate described by the inverse to the error exponent function. These results apply to both one-to-one and prefix-free variable rate codes.
In the case of universal compression, it is shown that the same performance can be universally achieved by one-to-one codes, at a cost of no more than $(\frac{m-2}{2})\log n$ bits, over the class of all memoryless sources on an alphabet of size $m$. The key technical step in the proof of this result is a precise evaluation of the number of strings with empirical entropy below a certain threshold.
There are several possible future directions in which the present results can be generalized and extended. Two immediate possibilities are described by the two conjectures we formulated in Section 1.4:
  1. 1.
    The universal rate obtained in Theorem 5.1 is optimal for one-to-one codes.
  2. 2.
    The best universally achievable rate for prefix-free codes incurs a penalty of $\frac{m}{2}\log n$ bits over the optimal (non-universal) rate. In other words, Theorem 5.1 remains valid for prefix-free codes, and a matching converse also holds, with $\frac{m-2}{2}$ replaced by $\frac{m}{2}$ in the universal rate $R^{\sf u}(n,\delta,P)$.
There are also several avenues for more ambitious generalizations. For example, it would be interesting to determine the corresponding pragmatic rate for the problem of lossy data compression or, even more interestingly, for channel coding. The results in this paper and the earlier hypothesis testing results discussed in the introduction are encouraging indications that it might be possible to identify similar phenomena in lossy data compression, in channel coding, and perhaps in some of the many other scenarios that arise in multi-terminal communication systems.